OpenStudy (anonymous):

Green's Teorem comprovate f(x,y)=y^3i+(x^3+3xy^2)j

OpenStudy (anonymous):

take gradient of f(x,y). so find fx and fy

OpenStudy (anonymous):

take q = x^3+3xy^2 and p = y^3

OpenStudy (anonymous):

dq/dx = 3x^2+3y^2, dp/dy = 3Y^2

OpenStudy (anonymous):

integral of dq/dx - dp/dy integral of (3x^2 + 3y^2 - 3y^2) integral of 3x^2

OpenStudy (anonymous):

x^3 + C

OpenStudy (anonymous):